A consumer’s demand curve for X is given by the equation P=100-squarerootofQ.

Calculate his point Price elasticity of demand when the price of Xi s 60.


1
Expert's answer
2021-03-31T07:25:59-0400

The price elasticity of demand can be found as follows:


"\epsilon=\dfrac{p}{q}\dfrac{dq}{dp}."

Let's write the direct demand function:


"\sqrt{q}=100-p,""q=p^2-200p+10000."

Then,


"\dfrac{dq}{dp}=2p-200."

Finally, we get:


"\epsilon=\dfrac{p}{p^2-200p+10000}(2p-200),""\epsilon=\dfrac{60}{60^2-200\cdot60+10000}\cdot(2\cdot60-200)=-3.0."